# Pisa, Italy

### v1

@@ Abstract
Index Terms:
Recurrent Neural Networks, Reservoir Computing, Echo State Networks, Time-series classification
@@ I Introduction
In this paper, we introduce Residual Reservoir Memory Networks (ResRMNs). ResRMNs are characterized by a hierarchical and modular structure, combining a linear reservoir and a non-linear reservoir. In particular, we consider an implementation for non-linear reservoir which is based on temporal residual connections. Through extensive empirical evaluation on time-series classification and pixel-level 1-D classification tasks, we demonstrate the advantages of the proposed approach over other RC-based models.
@@ III Residual Reservoir Memory Network
Similarly to [9], we will consider different ResRMN configurations based on the specific structure employed for the orthogonal matrix $\mathbf{O}$. More specifically, we will consider a random variant ResRMNR where matrix $\mathbf{O}$ is a random orthogonal matrix obtained via QR decomposition of a random matrix $\mathbf{M}^{N_{h}\times N_{h}}$ with i.i.d entries in (1,1)(-1,1), and two structured configurations ResRMNC and ResRMNI where the orthogonal matrix is implemented as a cyclic orthogonal matrix $C$, whose structure is highlighted in (4), and as the identity matrix $\mathbf{I}$, respectively.
@@ IV-A Eigenspectrum analysis
The different ResRMN configurations (random orthogonal, cyclic orthogonal, and identity) are intentionally selected so that we may have similar or dissimilar eigenspectra between the linear and non-linear modules. Specifically, similar eigenspectra are expected when the ResESN module employs either a cyclic orthogonal matrix, which matches the structure of the memory cell’s recurrent weight matrix $\mathbf{V_{m}}$, or a random orthogonal matrix, which exhibits a comparable eigenvalues’ distribution to that of the cyclic orthogonal matrix (see Fig. 2). Conversely, when the ResESN module employs the identity matrix, the modules should exhibit less redundancy due to the dissimilar eigenspectrum that characterizes the identity matrix. Therefore, conceptually, ResRMNI may capture a broader diversity of representations compared to the other configurations. In Fig. 3 we visualize the eigenspectrum of the Jacobian defined in (8) for each ResRMN configuration, differentiating between the $N_{m}$ eigenvalues of the memory module and the $N_{h}$ eigenvalues of the ResESN module. Trivially, the set of $N_{m}$ eigenvalues of the linear module are spread around the unitary circle due to its the recurrent weight matrix being initialized to a cyclic orthogonal structure. Conversely, the distribution of the $N_{h}$ eigenvalues of the non-linear module varies depending on which orthogonal matrix has been employed in the ResESN component. In particular, the random orthogonal and the cyclic orthogonal configurations are characterized by eigenvalues that spread around a $\beta\lVert\mathbf{W_{h}}\rVert$-neighborhood of the circle of radius $\alpha$. Conversely, the identity configuration is characterized by a more skewed eigenvalues’ distribution, where the eigenvalues fall in a $\beta\lVert\mathbf{W_{h}}\rVert$-neighborhood of $\alpha$. In the future it could be interesting to complement this analysis by exploring the eigenspectra in polar form, investigating the differences between each configuration by a different perspective that involves eigenvalues’ magnitudes and angles.
Figure 3: Eigenvalues of the Jacobian for different ResRMN configurations. The dynamics are driven by a random input vector and random states, both uniformly distributed in (1,1)(-1,1). We assume $N_{m},N_{h}=100$, a spectral radius $\rho=1$, all input weight matrices with scaling of 11 (i.e., $\omega_{x},\omega_{x_{m}},\omega_{m}=1$), zero bias $\omega_{b}=0$, and scaling coefficients $\alpha,\beta=1$. In red the $N_{m}$ eigenvalues of the linear memory module, in blue the $N_{h}$ eigenvalues of the ResESN module. In orange the unitary circle.
@@ V-A Model selection
V-B Results on time-series classification tasks
Table III presents the results for the considered classification tasks for a hidden size of $N_{h}=100$. To ensure a fair comparison, all models have the same number of trainable parameters. ResRMN demonstrates superior performance in the majority of datasets, particularly when taking into account the configuration employing the identity matrix in the temporal residual connections. Furthermore, the configuration that employs the cyclic orthogonal matrix in both modules is generally inferior even to the random orthogonal configuration. Fig. 4 (right) summarizes the average test accuracy improvements achieved by each class of models (ResESN, RMN, and ResRMN) compared to the leakyESN baseline across all time-series classification datasets. ResRMN, on average, outperforms the leakyESN by a $20.7\%$ gain in test accuracy, while also surpassing both ResESN and RMN. Finally, Fig. 4 (left) analyzes how the number of memory reservoir’s neurons $N_{m}$ affects performance in RMNs and ResRMNs. More specifically, we trained the models for values of $N_{m}$ in $\frac{1}{10}T,\frac{1}{5}T,T,2T]$, where $T$ is the sequence length of the dataset, computed performance changes relative to the base case when $N_{m}=T$, and averaged these changes across all datasets. We observe that $N_{m}$ has a crucial effect on performance, with all models decreasing in performance by at least $10\%$ for the lowest number of memory neurons considered (i.e., $N_{m}=\frac{1}{10}T$).
@@ V-B Results on time-series classification tasks
Figure 4: (a) Performance change in time-series classification tasks relative to the number of memory reservoir neurons $N_{m}$, with the base case being $N_{m}=T$. Results are broken down for non-residual and residual reservoir memory networks and averaged across all datasets. (b) Performance change in time-series classification tasks relative to leakyESN. Results are broken down for the different model classes considered in the experiments and averaged across all datasets.
@@ V-C Results on pixel-level classification tasks
Figure 5: Test results on psMNIST for an increasing number of trainable parameters. For RMN and ResRMNs, the number of memory reservoir neurons $N_{m}$ is fixed to $784$. Reported results represent the average, and the corresponding standard deviation, across $5$ random initializations.
@@ VI Conclusions
In the future, we plan to experiment with different initialization approaches for the memory cell’s recurrent weight matrix and more advanced architectures for the linear memory reservoir. Future work could also analyze the eigenvalues of the resulting models in terms of their polar form (magnitude and angle) rather than their cartesian form (real and imaginary parts). Leveraging this different perspective could provide insights into the differences between the orthogonal configurations and the identity one, particularly in understanding the rotational dynamics and stability properties introduced by each configuration.
Acknowledgment
Computational resources provided by Computing@Unipi, a computing service provided by University of Pisa.

### v2

# 1st Matteo Pinna    2nd Andrea Ceni    3rd Claudio Gallicchio

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/IJCNN64981.2025.11227859.
@@ I Introduction
In this paper, we introduce Residual Reservoir Memory Networks (ResRMNs). ResRMNs are characterized by a hierarchical and modular structure, combining a linear reservoir and a non-linear reservoir. Conceptually, the linear reservoir is designed to enhance memorization capabilities, while the non-linear reservoir, whose implementation in this work is based on temporal residual connections, is designed to capture complex, non-linear dependencies. We study different configurations for the temporal residual connections, including orthogonal matrices with either random or fixed structure. Additionally, we provide a characterization of the spectrum of ResRMN’s Jacobian in Th. 1, and analyze stability properties through the lens of linear stability analysis. Through extensive empirical evaluation on time-series classification and pixel-level 1-D classification tasks, we demonstrate the advantages of the proposed approach over other RC-based models.
@@ III Residual Reservoir Memory Network
Similarly to [9], we will consider different ResRMN configurations based on the specific structure employed for the orthogonal matrix $\mathbf{O}$. More specifically, we will consider a random variant ResRMNR where matrix $\mathbf{O}$ is a random orthogonal matrix obtained via QR decomposition of a random matrix $\mathbf{M}^{N_{h}\times N_{h}}$ with i.i.d entries in [1,1)[-1,1), and two structured configurations ResRMNC and ResRMNI where the orthogonal matrix is implemented as a cyclic orthogonal matrix $\mathbf{C}$, whose structure is highlighted in (4), and as the identity matrix $\mathbf{I}$, respectively. We choose these three configurations to cover a broad spectrum of dynamical behaviors [16]. In particular, the identity matrix is expected to provide relatively low memory capacity, while the other orthogonal configurations are expected to maximize it [17].
@@ IV Linear Stability Analysis
Figure 3: Eigenvalues of the Jacobian for different ResRMN configurations. The dynamics are driven by a random input vector and a random state, both uniformly distributed in (1,1)(-1,1). We assume hidden sizes $N_{m},N_{h}=100$, a spectral radius $\rho=1$, all input weight matrices with scaling of 11 (i.e., $\omega_{x},\omega_{x_{m}},\omega_{m}=1$), zero bias $\omega_{b}=0$, and scaling coefficients $\alpha,\beta=1$. In red the $N_{m}$ eigenvalues of the linear memory module, in blue the $N_{h}$ eigenvalues of the ResESN module. In orange the unitary circle.
@@ IV-A Eigenspectrum analysis
The different ResRMN configurations (random orthogonal, cyclic orthogonal, and identity) are intentionally selected so that we may have similar or dissimilar eigenspectra between the linear and non-linear modules. Specifically, similar eigenspectra are expected when the ResESN module employs either the cyclic orthogonal matrix, which matches the structure of the memory cell’s recurrent weight matrix $\mathbf{V}_{m}$, or a random orthogonal matrix, which exhibits a very similar eigenvalues’ distribution to that of the cyclic orthogonal matrix (see Fig. 2). Conversely, when the ResESN module employs the identity matrix, the modules should exhibit less redundancy in terms of their eigenspectra due to the dissimilar eigenspectrum that characterizes the identity matrix. Therefore, conceptually, ResRMNI may capture a broader diversity of representations compared to the other configurations. In Fig. 3 we visualize the eigenspectrum of the Jacobian defined in (8) for each ResRMN configuration, differentiating between the $N_{m}$ eigenvalues of the memory module and the $N_{h}$ eigenvalues of the non-linear module. Trivially, the set of $N_{m}$ eigenvalues of the linear module are spread around the unitary circle, due to its linear dynamics and its recurrent weight matrix being initialized to a cyclic orthogonal structure. Conversely, the distribution of the $N_{h}$ eigenvalues of the non-linear module varies depending on which orthogonal matrix is employed in the temporal residual connections. The random orthogonal and the cyclic orthogonal configurations are characterized by eigenvalues that spread around a $\beta\lVert\mathbf{W}_{h}\rVert$-neighborhood of the circle of radius $\alpha$. The identity configuration is characterized by a more skewed eigenvalues’ distribution, where the eigenvalues fall in a $\beta\lVert\mathbf{W}_{h}\rVert$-neighborhood of $\alpha$.
In future work, this analysis could be complemented by exploring the eigenspectra in polar form, investigating the differences between each configuration from the perspective of eigenvalues’ magnitudes and angles, rather than their real and imaginary parts.
@@ V Experiments
Figure 4: (a) Performance change in time-series classification tasks relative to the number of memory reservoir neurons $N_{m}$, with the base case being $N_{m}=T$. Results are broken down for non-residual and residual reservoir memory networks and averaged across all datasets. (b) Performance change in time-series classification tasks relative to LeakyESN. Results are broken down by model class and averaged across all datasets.
V-A Model selection
Figure 5: Test results on psMNIST for a varying number of trainable parameters. For RMN and ResRMNs, the number of memory reservoir neurons $N_{m}$ is fixed to $784$. Reported results represent the average, and the corresponding standard deviation, across $5$ random initializations.
V-B Results on time-series classification tasks
Table III presents the results for the considered classification tasks for a hidden size of $N_{h}=100$. To ensure a fair comparison, all models have the same number of trainable parameters. ResRMN demonstrates superior performance in the majority of datasets, particularly when taking into account the configuration employing the identity matrix in the temporal residual connections. Furthermore, the configuration that employs the cyclic orthogonal matrix in both modules is generally inferior even to the random orthogonal configuration. In some datasets, we observe high degrees of variability (i.e., standard deviations). We argue that this variability may be related to the intrinsic characteristics of each specific dataset. In future work, we plan to investigate in detail how this variability changes as the number of units in the reservoir increases.
Fig. 4 (right) summarizes the average test accuracy improvements achieved by each class of models (ResESN, RMN, and ResRMN) compared to the LeakyESN baseline across all time-series classification datasets. ResRMN, on average, outperforms the LeakyESN by a $19.9\%$ gain in test accuracy, while also surpassing both ResESN and RMN. Finally, Fig. 4 (left) analyzes how the number of memory reservoir’s neurons $N_{m}$ affects performance in RMNs and ResRMNs. More specifically, we trained these models for values of $N_{m}$ in $\frac{1}{10}T,\frac{1}{5}T,T,2T]$, where $T$ is the sequence length of the dataset, computed performance changes relative to the base case when $N_{m}=T$, and averaged these changes across all datasets. We observe that $N_{m}$ has a crucial effect on performance, with all modular models decreasing in performance by at least $10\%$ for the lowest number of memory neurons considered (i.e., $N_{m}=\frac{1}{10}T$). In the future, we plan to analyze, with respect to $N_{m}$, the trade-off between performance and training time to investigate the overhead introduced by leveraging two reservoirs rather than one.
@@ VI Conclusions
In future work, we plan to experiment with different initialization approaches for the memory cell’s recurrent weight matrix and more advanced architectures for the linear memory reservoir. Future work could also analyze the eigenvalues of the proposed model in terms of their polar form (magnitude and angle) rather than their cartesian form (real and imaginary parts). Leveraging this different perspective could provide insights into the differences between the orthogonal configurations and the identity one, particularly in understanding the rotational dynamics and stability properties introduced by each configuration.
Acknowledgments
Computational resources were provided by Computing@Unipi, a computing service of the University of Pisa.